# Feature Calibration Enhanced Parameter Synthesis for CLIP-based Class-Incremental Learning

### v4

@@ Abstract
Class-Incremental Learning (CIL) enables models to continuously learn new class knowledge while retaining previous classes, facilitating adaptation and evolution in dynamic, real-world environments. Traditional CIL methods primarily rely on visual features, which limits their effectiveness in complex, multimodal scenarios. In contrast, VLMs show promising potential for enhancing CIL by leveraging pre-trained knowledge and integrating multi-modal semantic cues such as text and vision. However, existing approaches struggle to mitigate catastrophic forgetting while preserving the generalization strengths of VLMs across diverse modalities. To address these challenges, we propose a Feature Calibration Enhanced Parameter Synthesis (FCPS) framework. Specifically, FCPS introduces a dynamic parameter adjustment mechanism that iteratively calibrates the contribution of original visual features to the final class decision, thus preserving the model’s intrinsic generalization capability across modalities. Simultaneously, parameter integration enables effective knowledge transfer, maintaining a balance between acquiring new class representations and preserving old knowledge. Experimental results on popular benchmarks (e.g., CIFAR100 and ImageNet100) validate the superiority of the proposed method.
Index Terms:
@@
Class-Incremental Learning, Visual Language Learning, Knowledge Distillation, Feature Calibration.
I Introduction
@@
Class-Incremental Learning (CIL) aims to incorporate the knowledge of new classes incrementally and build a universal classification model among all seen classes [1, 2, 3, 4, 5]. In other words, CIL should not only continuously acquire knowledge from the current task but also preserve the knowledge from former tasks, making it crucial for real-world applications [6, 7] that require adaptability. For example, autonomous driving systems must dynamically adapt to evolving road conditions, weather variations, traffic regulations, and driving patterns [8], while medical imaging systems face challenges from emerging diseases and evolving pathological presentations. These scenarios require CIL-enabled models that effectively integrate new information without revisiting previous data - a necessity for deployment on edge devices constrained by limited storage and computational capacity.
Figure 1: Illustration of existing CIL framework.
@@
A primary challenge in CIL is to mitigate catastrophic forgetting [9], where the integration of new knowledge could overwrite previously acquired information in the process of continuously building a classifier for all classes. Effective solutions must balance stability—the retention of past knowledge—and plasticity—the adaptability to new tasks [10]. To address this, various approaches have been proposed, including replay-based[11], regularization-based [12], and parameter-isolated methods [13]. For instance, [11] proposes an efficient memory scheme based on the Resnet-18 to cope with the highly imbalanced data problem, while DNE [14] introduces a dynamic architecture that balances accuracy and model complexity based on Vision Transformer (ViT)[15] backbone. While these strategies have shown success with foundational vision backbones (as shown in Fig. 1(a)) like ResNet[16] and ViT[15], they focus solely on visual features, limiting their robustness in complex, multi-modal scenarios where integrating heterogeneous information sources is essential.
@@
Recent advancements in incremental learning have explored Vision-Language Models (VLMs), such as ViT-BERT [17] and CLIP [18], which harness both visual and textual information to enhance performance. These pretrained models have learned rich feature representation from large-scale multi-modal data [19, 20], which can be used to bridge the knowledge gap between class information trained on different stages. By freezing CLIP’s pretrained backbone and leveraging its zero-shot capabilities, Continual-CLIP[21] demonstrates the pretained VLMs are well suited for CIL. Motivated by the zero-shot ability of CLIP, the follow-up methods propose different finetuning methods to mitigate catastrophic forgetting, consistently outperforming traditional vision-only methods [22].
@@
Prompt-tuning[23, 24] and visual feature tuning [25] are two representative categories of CLIP-based approaches for incremental learning. Prompt-tuning employs task-specific or input-conditioned prompts to guide the model while preserving the original backbone, while visual feature tuning modifies the pretrained model to accommodate new tasks as illustrated in Fig. 1(c). Even though the previous methods can inject valuable pretrained knowledge, they still face challenges for CIL. First, prompt-tuning methods primarily focus on improving and enhancing the text prompt space in Fig. 1(b), overlooking the importance of effectively utilizing visual features. Second, visual feature tuning methods risk degrading its feature extraction ability based on its coverage of the original CLIP features. As shown in Fig. 2, the last-stage classification clusters of visual feature tuning method is not even seperated clearly compared to those of zero-shot CLIP. The recently proposed RAPF[26] is a visual feature tuning method that uses neighboring categories separation to enhance the representation of the feature space. However, the new and old knowledge is projected into one shared feature basis, limiting the ability to adapt to the shared feature space dynamically and adopting the CLIP features.
@@
To address these issues, this paper proposes a novel CIL method called Feature Calibration Enhanced Parameter Synthesis (FCPS). We adopt the frozen encoder of CLIP as the foundational architecture for our approach. Considering the potential distortion of the pretrained knowledge of CLIP caused by fine-tuning, we designed Feature Calibration (FC) which uses the trainable layer to transform the CLIP features into new feature spaces. FC effectively utilizes and rationalizes the enhancement of the original visual features of CLIP while introducing specialized parameters to control the retention of the original visual features of CLIP precisely. It allows the model to maintain strong generalization capabilities while reducing the likelihood of overfitting. At the same time, we focus on whether the model can be dynamically adapted between old and new knowledge in the process of continuously building a classifier for all classes, rather than on the adaptation and improvement of textual cues. We introduce the Parameter Synthesis (PS) module, which effectively mitigates catastrophic forgetting by addressing the significant discrepancies in knowledge between adjacent tasks. PS integrates the adapter parameters of the current stage with those of the previous stage to initialize the adapter parameters for the next stage. With task evolution, the parameters of adjacent stages undergo decomposition and integration, alleviating the forgetting issue as tasks progress. Furthermore, we incorporate an additional dynamic distillation loss during the training process to constrain the changes of adapter parameters, further mitigating forgetting. The main contributions of this paper are as follows:
  • We propose a tuning-efficient VLM-based framework for CIL, which effectively solves the issue of the original feature of CLIP being distorted due to fine-tuning with a feature calibration module.
  • To tackle the catastrophic forgetting problem, we introduce the QR decomposition-based parameter synthesis, which captures parameter changes during task transitions and fuses the parameters constructed for different tasks.
@@
The remainder of this paper is organized as follows. In Section II, we review related work on class-incremental learning and Vision-Language Models. Section III introduces the overall framework of our proposed method. We then elaborate on the design of the feature calibration and parameter synthesis modules, and clarify the optimization objective. In Section IV, we present comprehensive experimental results and compare FCPS with state-of-the-art methods. Finally, Section V concludes the paper by discussing limitations and potential future research directions.
II Related work
II-A Class-Incremental Learning
@@
CIL aims to continuously incorporate new knowledge from new data while retaining, integrating and even optimizing old knowledge to prevent catastrophic forgetting [10] in the process of continuously building a classifier for all classes [27]. The current mainstream methods for CIL are divided into three main categories. The replay-based methods either save a portion of samples from previous tasks [28, 29, 30, 31] or generate synthetic samples with a generative model [32, 33, 34, 35, 36, 37, 38] in a cache. When training on new tasks, these samples are used alongside new samples to prevent forgetting previously learned knowledge. [29] selects samples with high predictive entropy and proximity to the decision boundary, aiming to enhance generalization by replaying these challenging examples. [28] estimates sample uncertainty through data augmentation, identifying instances with high prediction variance across multiple augmented versions. Regularization-based methods impose constraints on network parameter updates. Weight regularization focuses on preventing drift in weights associated with previous tasks [39, 12, 40] , while data regularization methods transfer knowledge from the old model to the new one through knowledge distillation [41, 42, 43] . Given that each parameter does not contribute equally to the task, parameter regularization methods seek to assess the importance of each parameter to the network and keep important parameters static to preserve previous knowledge. [40] examines the performance and stability of various network layers in CIL, revealing distinct characteristics. Shallow layers learn quickly but offer limited representational capacity, whereas deeper layers adapt more slowly yet provide stronger discriminative features. While knowledge distillation refers to building a mapping between a set of methods, which has been widely adopted in CIL in several ways. AFC [42] considers the importance of different feature maps for distillation. COIL [41] suggests the use of collaborative transport for bidirectional distillation, where both the semantic relationships between old and new models are exploited. Parameter isolation-based methods mitigate catastrophic forgetting by dynamically increasing the capacity of network. It enables the model to retain knowledge from previous tasks while expanding to learn new ones [44, 13, 45, 46, 47]. [46] proposes the Dynamically Expandable Representation (DER) method, which decouples feature adaptation from the classifier. For a new task, DER freezes the existing feature extractor, adds a new one, and concatenates features from all extractors for classification. [47] proposes a dynamic structure reorganization strategy that preserves old knowledge while learning new classes. When a new class arrives, side branches are added to the network blocks. The original weights are frozen, and the new branches are trained to learn the new class. Due to the excellent zero-shot capability of CLIP on single or multi-task [48, 49, 50, 51, 20], recent works [25, 23, 24] have shown the possibility of applying CLIP to incremental learning scenarios.
II-B Vision-Language Models
@@
Large-scale Vision-Language Models such as CLIP [18], ALIGN [52], and Florence [53] have been pretrained on massive datasets containing image-text pairs, enabling them to learn a unified embedding space that bridges the gap between visual and textual modalities [54]. These models exhibit remarkable generalization and transfer capabilities, making them particularly effective for zero-shot learning across a wide range of downstream tasks [55] without the need for task-specific fine-tuning. In the context of incremental learning, a growing body of research has begun to explore how to adapt these powerful pretrained models to sequential tasks while minimizing catastrophic forgetting [56]. A recent trend involves introducing additional learnable components—such as prompts or lightweight modules—that allow the pretrained backbone to remain largely unchanged while adapting to new tasks. For instance, L2P [57] brings prompt-based learning into incremental learning by maintaining a shared pool of prompts that guide the model to adapt to each task dynamically. DualPrompt [22] proposes attaching complementary prompts to the pretrained backbone and formulates the objective as learning task-invariant and task-specific instructions. Most of these works focus on prompt adjustment, prefix adjustment and effective organization of text features but ignore the effective attention to visual features. This paper aims to fill this research gap by proposing a comprehensive solution that focuses on visual features to tune visual language models without suffering from forgetting.
III Methodology
III-A Overview
@@
Traditional visual methods often struggle to make effective use of textual information [58], and fine-tuning methods based on CLIP have shown promising performance in multimodal settings. They are highly susceptible to catastrophic forgetting when applied to incremental learning scenarios. To address these limitations, we propose a novel method called FCPS, which effectively leverages the rich cross-modal knowledge encapsulated in CLIP while simultaneously addressing the challenge of forgetting. As shown in Fig.3, FCPS enhances performance through three core components that collectively aim to balance knowledge retention and task comprehensiveness. The first component FC is introduced to preserve and optimize the pretrained CLIP representations. To reduce the risk of distorting the pretrained CLIP knowledge caused by direct fine-tuning, FC is designed to refine the extracted visual features in a lightweight and stable manner. This module enhances the visual embeddings without altering the core CLIP encoder, ensuring that the foundational multimodal knowledge is preserved. To further maintain the generalization capability of the model while preventing overfitting to the current task, we introduce a parameter $\alpha$ after the calibration stage. This parameter determines the extent to which the original image embeddings from the CLIP Image Encoder are retained. By maintaining a proportion of these original embeddings, the model can better preserve the semantics of the pretrained representation space, leading to improved robustness across diverse tasks. The second core component, PS, is designed to address the issue of forgetting. It effectively integrates the current adapter parameters with those from the previous stage, serving as the initial parameters for the next stage’s adapter. This progressive update mechanism allows knowledge from earlier tasks to be preserved and integrated into future learning. As tasks evolve, PS accumulates experience by continuously adapting and refining its parameters, ensuring that knowledge from old classes is retained even as new ones are introduced. This design effectively mitigates catastrophic forgetting in the process of continuously building a classifier for all classes. Finally, FCPS introduces Dynamic Distillation, a module responsible for transferring knowledge between the current and previous tasks. This component acts as a bridge between stages, enhancing continuity and semantic consistency throughout the learning process. In summary, FCPS offers a comprehensive and systematic approach to incremental learning with VLMs. By preserving cross-modal knowledge, promoting effective parameter reuse, and introducing a distillation mechanism, FCPS achieves a strong balance between performance on new tasks and retention of previously acquired knowledge.
@@
Figure 3: The framework of the proposed feature calibration enhanced parameter synthesis.
III-B Feature Calibration
@@
III-C Parameter Synthesis
To address the challenge of catastrophic forgetting in incremental learning, we introduce parameter synthesis—a mechanism that dynamically integrates adapter parameters from the current learning stage with those preserved from previous stages. This integration enables the progressive fusion of new and old knowledge, ensuring that the model can incrementally construct a unified classifier capable of handling all previously encountered classes. By retaining and reusing historical parameters, PS promotes long-term memory retention while simultaneously adapting to novel information, thereby mitigating the risk of performance degradation due to forgetting.
In many traditional weight update schemes, newly acquired information often overwrites past knowledge, leading to a well-known issue of catastrophic forgetting. This problem is particularly pronounced in incremental learning scenarios, where the model is required to adapt continuously to a sequence of tasks without access to the full dataset from earlier stages. To further combat this issue, we propose a strategy based on QR decomposition. Specifically, during each incremental update, the model’s weight matrix is decomposed into an orthogonal matrix $\mathbf{Q}$ and an upper triangular matrix $\mathbf{R}$, allowing us to restructure and recombine the parameter space in a manner that preserves essential features of previously learned tasks. The use of QR decomposition offers a mathematically grounded approach to separating and retaining knowledge components from earlier stages, while facilitating the smooth integration of new task-specific information. By recombining these components in a controlled way, the model can maintain a balanced representation across tasks, effectively reducing interference and preserving critical knowledge.

### v5

# CalFuse: Feature Calibration Enhanced Parameter Fusion for Class-Continual Learning

@@ Abstract
Class-Continual Learning (CCL) enables models to continuously learn new class knowledge while retaining previous classes, facilitating adaptation and evolution in dynamic, real-world environments. Traditional CCL methods primarily rely on visual features, which limits their effectiveness in complex, multimodal scenarios. In contrast, Vision-Language Models (VLMs) show promising potential for enhancing CCL by leveraging pre-trained knowledge and fusing multi-modal semantic cues such as text and vision. However, existing approaches struggle to mitigate catastrophic forgetting while preserving the generalization strengths of VLMs across diverse modalities. To address these challenges, we propose CalFuse, a framework for feature Calibration enhanced parameter Fusion, which enhances dynamic knowledge fusion. CalFuse introduces a dynamic feature calibration mechanism that iteratively adjusts the contribution of original visual features to the final class decision, thereby preserving the model’s intrinsic generalization capability across modalities. Simultaneously, a parameter fusion strategy effectively fuses newly acquired knowledge with prior task parameters, maintaining a balance between acquiring new class representations and preserving old knowledge. Experimental results on popular benchmarks (e.g., CIFAR100 and ImageNet100) validate the superiority of the proposed method.
keywords:
@@
Class-Continual Learning, Visual Language Learning, Parameter Fusion, Knowledge Distillation
[…]
[1]organization=Fudan University,city=ShangHai, postcode=200433, country=China
[…]
[2]organization=University of California, Davis,postcode=95616, state=CA, country=USA
[…]
[3]organization=University of Rochester,postcode=14627, state=New York, country=USA
[…]
[4]organization=Ningbo University,postcode=315211, city=Ningbo, stateZhejiang Province country=China
[…]
[6]organization=Beijing University of Posts and Telecommunications,postcode=100193, state=BeiJing, country=China
[…]
[5]organization=Cardiff University,postcode=CF24 4AG, state=Wales, countr=UK
1 Introduction
@@
Class-Continual Learning (CCL) aims to incorporate the knowledge of new classes incrementally and build a universal classification model among all seen classes. In other words, CCL should not only continuously acquire knowledge from the current task but also preserve the knowledge from former tasks [1, 2, 3], which is crucial for real-world applications [4, 5, 6, 7, 8] that require adaptability. For example, autonomous driving systems must dynamically adapt to evolving road conditions, weather variations, traffic regulations, and driving patterns [9], while medical imaging systems face challenges from emerging diseases and evolving pathological presentations. These scenarios require CCL-enabled models that effectively integrate new information without revisiting previous data - a necessity for deployment on edge devices constrained by limited storage and computational capacity.
Figure 1: Illustration of existing CCL framework.
@@
A primary challenge in CCL is to mitigate catastrophic forgetting [10],where the integration of new knowledge could overwrite previously acquired information in the process of continuously building a classifier for all classes. Effective solutions must balance stability—the retention of past knowledge—and plasticity—the adaptability to new tasks [10]. To address this, various approaches have been proposed, including replay-based [11], regularization-based [12], and parameter-isolated methods [13]. For instance, [11] proposes an efficient memory scheme based on the Resnet-18 to cope with the highly imbalanced data problem, while DNE [14] introduces a dynamic architecture that balances accuracy and model complexity based on Vision Transformer (ViT) [15] backbone. While these strategies have shown success with foundational vision backbones (as shown in Fig. 1(a)) like ResNet [16] and ViT [15], they focus solely on visual features, limiting their robustness in complex, multi-modal scenarios where integrating heterogeneous information sources is essential.
@@
Recent advancements in continual learning have explored Vision-Language Models (VLMs), such as ViT-BERT [17] and CLIP [18], which harness both visual and textual information to enhance performance. These pretrained models have learned rich feature representation from large-scale multi-modal data , which can be used to bridge the knowledge gap between class information trained on different stages. By freezing CLIP’s pretrained backbone and leveraging its zero-shot capabilities, Continual-CLIP [19] demonstrates the pretained VLMs are well suited for CCL. Motivated by the zero-shot ability of CLIP, the follow-up methods propose different finetuning methods to mitigate catastrophic forgetting, consistently outperforming traditional vision-only methods [20].
@@
Prompt-tuning [21, 22] and visual feature tuning [23] are two representative categories of CLIP-based approaches for continual learning. Prompt-tuning employs task-specific or input-conditioned prompts to guide the model while preserving the original backbone, while visual feature tuning modifies the pretrained model to accommodate new tasks as illustrated in Fig. 1(c). Even though the previous methods can inject valuable pretrained knowledge, they still face challenges for CCL. First, prompt-tuning methods primarily focus on improving and enhancing the text prompt space in Fig. 1(b), overlooking the importance of effectively utilizing visual features. Second, visual feature tuning methods risk degrading its feature extraction ability based on its coverage of the original CLIP features. As shown in Fig. 2, the last-stage classification clusters of visual feature tuning method is not even seperated clearly compared to those of zero-shot CLIP.
@@
To address these issues, this paper proposes a novel CCL method called CalFuse, integrating dual mechanisms of Feature Calibration and enhanced Parameter Fusion. We adopt the frozen encoder of CLIP as the foundational architecture for our approach. Considering the potential distortion of the pretrained knowledge of CLIP caused by fine-tuning, we designed a Feature Calibration (FC) mechanism which uses the trainable layer to fuse original CLIP features into new feature spaces. FC not only enhances CLIP’s visual features by rationalizing their discriminative power but also introduces specialized parameters to precisely control the fusion ratio of original and adapted features, ensuring the model maintains strong cross-modal generalization while reducing overfitting. Concurrently, as opposed to solely focusing on textual cue adaptation, we prioritize the dynamic knowledge fusion across all classes between sequential tasks during classifier construction. We introduce the Parameter Fusion (PF) module, which effectively mitigates catastrophic forgetting by addressing the significant discrepancies in knowledge between adjacent tasks. PF fuses the adapter parameters of the current stage with those of the previous stage to initialize the adapter parameters for the next stage. With task evolution, the parameters of adjacent stages undergo decomposition and fusion, alleviating the forgetting issue as tasks progress. Furthermore, we incorporate an additional dynamic distillation loss during the training process to constrain the changes of adapter parameters, further mitigating forgetting. The main contributions of this paper are as follows:
  • 1.
    We propose a tuning-efficient VLM-based framework for CCL, which effectively solves the issue of the original feature of CLIP being distorted due to fine-tuning with a feature calibration module.
  • 2.
    To tackle the catastrophic forgetting problem, we introduce the QR decomposition-based parameter fusion, which captures parameter changes during task transitions and fuses the parameters constructed for different tasks.
  • 3.
@@
The remainder of this paper is organized as follows. In Section 2, we review related work on Class-Continual Learning and Vision-Language Models. Section 3 introduces the overall framework of our proposed method. We then elaborate on the design of the feature calibration and parameter fusion modules, and clarify the optimization objective. In Section 4, we present comprehensive experimental results and compare CalFuse with state-of-the-art methods. Finally, we conclude the paper by discussing limitations and potential future research directions.
2 Related work
2.1 Class-Continual Learning
@@
CCL aims to continuously incorporate new knowledge from new data while retaining, integrating and even optimizing old knowledge to prevent catastrophic forgetting [10] in the process of continuously building a classifier for all classes. The current mainstream methods for CCL are divided into three main categories. The replay-based methods either save a portion of samples from previous tasks or generate synthetic samples with a generative model in a cache [11]. When training on new tasks, these samples are used alongside new samples to prevent forgetting previously learned knowledge. [24] selects samples with high predictive entropy and proximity to the decision boundary, aiming to enhance generalization by replaying these challenging examples. [25] estimates sample uncertainty through data augmentation, identifying instances with high prediction variance across multiple augmented versions. Regularization-based methods impose constraints on network parameter updates. Weight regularization focuses on preventing drift in weights associated with previous tasks [12], while data regularization methods transfer knowledge from the old model to the new one through knowledge distillation [26]. Given that each parameter does not contribute equally to the task, parameter regularization methods seek to assess the importance of each parameter to the network and keep important parameters static to preserve previous knowledge. [27] examines the performance and stability of various network layers in CCL, revealing distinct characteristics. Shallow layers learn quickly but offer limited representational capacity, whereas deeper layers adapt more slowly yet provide stronger discriminative features. Knowledge distillation refers to building a mapping between a set of methods, which has been widely adopted in CCL in several ways. COIL [26] suggests the use of collaborative transport for bidirectional distillation, where both the semantic relationships between old and new models are exploited. Parameter isolation-based methods mitigate catastrophic forgetting by dynamically increasing the capacity of network. It enables the model to retain knowledge from previous tasks while expanding to learn new ones [13]. [13] proposes the Dynamically Expandable Representation (DER) method, which decouples feature adaptation from the classifier. For a new task, DER freezes the existing feature extractor, adds a new one, and concatenates features from all extractors for classification. [28] proposes a dynamic structure reorganization strategy that preserves old knowledge while learning new classes. When a new class arrives, side branches are added to the network blocks. The original weights are frozen, and the new branches are trained to learn the new class. Due to the excellent zero-shot capability of CLIP on single or multi-task [29], recent works [23, 21, 22] have shown the possibility of applying CLIP to continual learning scenarios.
2.2 Vision-Language Models
@@
Large-scale Vision-Language Models such as CLIP [18], ALIGN [30], and Florence [31] have been pretrained on massive datasets containing image-text pairs, enabling them to learn a unified embedding space that bridges the gap between visual and textual modalities. These models exhibit remarkable generalization and transfer capabilities, making them particularly effective for zero-shot learning across a wide range of downstream tasks without the need for task-specific fine-tuning. In the context of continual learning, a growing body of research has begun to explore how to adapt these powerful pretrained models to sequential tasks while minimizing catastrophic forgetting [32]. A recent trend involves introducing additional learnable components—such as prompts or lightweight modules—that allow the pretrained backbone to remain largely unchanged while adapting to new tasks. For instance, L2P [33] brings prompt-based learning into continual learning by maintaining a shared pool of prompts that guide the model to adapt to each task dynamically. DualPrompt [20] proposes attaching complementary prompts to the pretrained backbone and formulates the objective as learning task-invariant and task-specific instructions. Most of these works focus on prompt adjustment, prefix adjustment and effective organization of text features but ignore the effective attention to visual features. This paper aims to fill this research gap by proposing a comprehensive solution that focuses on visual features to tune visual language models without suffering from forgetting.
3 Methodology
3.1 Overview
@@
Traditional visual methods often struggle to make effective use of textual information, and fine-tuning methods based on CLIP have shown promising performance in multimodal settings. They are highly susceptible to catastrophic forgetting when applied to continual learning scenarios. To address these limitations, we propose a novel method called CalFuse, which effectively leverages the rich cross-modal knowledge encapsulated in CLIP while simultaneously addressing the challenge of forgetting. As shown in Fig.3, CalFuse enhances performance through three core components that collectively aim to balance knowledge retention and task comprehensiveness. The first component FC is introduced to preserve and optimize the pretrained CLIP representations. To reduce the risk of distorting the pretrained CLIP knowledge caused by direct fine-tuning, FC is designed to refine the extracted visual features in a lightweight and stable manner. This module enhances the visual embeddings without altering the core CLIP encoder, ensuring that the foundational multimodal knowledge is preserved. To further maintain the generalization capability of the model while preventing overfitting to the current task, we introduce a parameter $\alpha$ after the calibration stage. This parameter determines the extent to which the original image embeddings from the CLIP Image Encoder are retained. By maintaining a proportion of these original embeddings, the model can better preserve the semantics of the pretrained representation space, leading to improved robustness across diverse tasks. The second core component, PF, is designed to address the issue of forgetting. It effectively fuses the current adapter parameters with those from the previous stage, serving as the initial parameters for the next stage’s adapter. This progressive update mechanism allows knowledge from earlier tasks to be preserved and fused into future learning. As tasks evolve, PF accumulates experience by continuously adapting and refining its parameters, ensuring that knowledge from old classes is retained even as new ones are introduced. This design effectively mitigates catastrophic forgetting in the process of continuously building a classifier for all classes. Finally, CalFuse introduces Dynamic Distillation, a module responsible for transferring knowledge between the current and previous tasks. This component acts as a bridge between stages, enhancing continuity and semantic consistency throughout the learning process. In summary, CalFuse offers a comprehensive and systematic approach to continual learning with VLMs. By preserving cross-modal knowledge, promoting effective parameter reuse, and introducing a distillation mechanism, CalFuse achieves a strong balance between performance on new tasks and retention of previously acquired knowledge.
@@
Figure 3: The framework of the proposed feature calibration enhanced parameter fusion.
3.2 Feature Calibration
@@
3.3 Parameter Fusion
To address the challenge of catastrophic forgetting in continual learning, we introduce parameter fusion—a mechanism that dynamically fuses adapter parameters from the current learning stage with those preserved from previous stages. It enables the progressive fusion of new and old knowledge, ensuring that the model can incrementally construct a unified classifier capable of handling all previously encountered classes. By retaining and reusing historical parameters, PF promotes long-term memory retention while simultaneously adapting to novel information, thereby mitigating the risk of performance degradation due to forgetting.
In many traditional weight update schemes, newly acquired information often overwrites past knowledge, leading to a well-known issue of catastrophic forgetting. This problem is particularly pronounced in continual learning scenarios, where the model is required to adapt continuously to a sequence of tasks without access to the full dataset from earlier stages. To further combat this issue, we propose a strategy based on QR decomposition. Specifically, during each incremental update, the model’s weight matrix is decomposed into an orthogonal matrix $\mathbf{Q}$ and an upper triangular matrix $\mathbf{R}$, allowing us to restructure and recombine the parameter space in a manner that preserves essential features of previously learned tasks. The use of QR decomposition offers a mathematically grounded approach to separating and retaining knowledge components from earlier stages, while facilitating the smooth integration of new task-specific information. By recombining these components in a controlled way, the model can maintain a balanced representation across tasks, effectively reducing interference and preserving critical knowledge.